if the ratio of minor disasters to major disasters was 5 to 2 and there were 980 disasters in all how many were minor
Answer by jim_thompson5910(35256)

Let

x = # of minor disasters
y = # of major disasters

5/2 = x/y

5y = 2x

2x = 5y

x = 5y/2

-------------------------------------------------------

x+y = 980

5y/2 + y = 980

2*(5y/2 + y) = 2*980

5y + 2y = 1960

7y = 1960

y = 1960/7

y = 280

x = 5y/2

x = 5*280/2

x = 700

So there were 700 minor disasters.